OpenStudy (anonymous):

Find all real solutions of 1/x^2 + 11/x +28 = 0

OpenStudy (timaashorty):

\[\frac{ 1 }{ x^2 } + \frac{ 11 }{ x }+28 = 0 \] like that?

OpenStudy (timaashorty):

If so, \[x= -\frac{ 1 }{ 4} , \ -\frac{ 1 }{ 7}\]

OpenStudy (timaashorty):

Those are the solutions

OpenStudy (anonymous):

Yes

OpenStudy (timaashorty):

you're welcome lol

OpenStudy (anonymous):

How did you come up with those answers?